OpenStudy (anonymous):

Hi everyone! Please explain how to find the limit as "t" goes to infinity of (-sin(2t)) / (se^(st))...this looks like an improper/improper which needs L'ophital's rule but not sure how to implement that. Anyone? Thanks! :o)

OpenStudy (anonymous):

now that I think about it, if sin(2t) bounces back and forth forever between y=-1 and y=1, then the numerator is a real number and the denominator goes to infinity, which would make the lim go to zero correct?

OpenStudy (rational):

thats right only if s > 0

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

well how would I know if s>0?

OpenStudy (rational):

you very much seem to be doing laplace transforms ?

OpenStudy (anonymous):

lol...yes! :o)

OpenStudy (rational):

then s > 0 is assumed

OpenStudy (anonymous):

this may seem like a dumb question but, why is it assumed?

OpenStudy (rational):

you wont ask that question if you look at the integral that you use to find laplace transform

OpenStudy (anonymous):

not sure I understand what you mean

OpenStudy (rational):

look at the definition of laplace transform in your notes

OpenStudy (rational):

and ask yourself for what values of "s" the integral makes sense (converges)

OpenStudy (anonymous):

oh yes....I remember...integral is between zero and infinity, so it's assumed that s>0

OpenStudy (anonymous):

actually, that wasn't a dumb question, it helped me remember the root of what I'm doing!

OpenStudy (rational):

it is a good question, i didnt answer directly because i want you look at that laplace transform integral haha

OpenStudy (anonymous):

if s is a negative, then the LaPlace definition diverges! :O)

OpenStudy (rational):

as simple as that

OpenStudy (anonymous):

yay!...now back to my original problem...this was only a portion of one of those nasty ones where you have to set your work equal to the original laplace transform and do all the tricky cancelling! bleh! :o/

OpenStudy (anonymous):

Thanks rational!

OpenStudy (rational):

you're doing good, yes that limit evaluates to 0

OpenStudy (rational):

yw and good luck!

OpenStudy (rational):

If you need to justify well, you may use squeeze thm

OpenStudy (rational):

\[-1\le\sin(2t)\le1\]

OpenStudy (anonymous):

i don't even remember squeeze theorem...but not sure I can take the time to relearn right now...maybe later

OpenStudy (rational):

kk squeeze thm is for real analysis, rigor is not needed when you're in diff eqns ;p

OpenStudy (anonymous):

good! whew!

OpenStudy (anonymous):

I will be doing math for the next 7 hours...will you be around?

OpenStudy (rational):

Yes feel free to tag me :)

OpenStudy (anonymous):

thanks @rational

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

quick question however... do you know what those weird squiggly "u's" are for?

OpenStudy (rational):

unit step functions ?

OpenStudy (anonymous):

something about step functions maybe?

OpenStudy (anonymous):

yes I think so

OpenStudy (anonymous):

can you explain those at all...like what is the purpose of the symbol?

OpenStudy (rational):

it is a piecewise function : \[\large u_c(t)=\left\{\begin{array}{}0&:&t\lt c\\1&:&t\ge c\end{array}\right.\]

OpenStudy (rational):

for example, \(\large u_0(t)\) looks like this :